Check my work Problem 17-1 ABO calculations; present value concepts [LO17-2, 17-3) Sachs Brands defined benefit pension plan specifies annual retirement benefits equal to: 1.6% service payable at the end of each year Angela Davenport was hired by Sachs at the beginning of 2004 a 2038 after 35 years service. He companys actuary projects her salary to be $325,000 at retirement. The actuarys discount rate is 9%. EvoSL Pyosi PVA of $1. EVAD of $1 and PVAD of S) (Use appropriate factor(s) from the tables provided.) years final years salary. r retirement is expected to span 18 years. Davenports salary is $99,000 at the end of 2018 and the . EAS. Required: 2. Estimate by the accumulated benefits approach the amount of Davenports annual retirement payments earned as of the endof 2018. 3. What is the companys accumulated benefit obligation at the end of 2018 with respect to Davenport? (Do not round intermediate calculations. Round your final answer to nearest whole dollar.) 4. If no estimates are changed in the meantime, what will be the accumulated benefit obligation at the end of 2021 (three years later) when Davenports salary is $160,000? (Do notFound intermediate calculations. Round your final answer to nearest whole dollar.) Et 2. Annual retirement payments 3. Accumulated benefit obligation 2018 4. Accumulated benefit obligation 2021 KPrey 11 of ext
Answer #1

2. Th amount of Davenport's annual retirement payments earned as of the end of 2013 :-

1.6 % * 15 * $ 99,000 = $ 23760

3. Th present value of the annuity as of the retirement date is :-

$ 23,760 * PVIFA ( 9%,18) = $23,760 * 8.7556 = $ 208,033.

The ABO is the present value of the retirement benefits at the end of 2013:-

$ 208,033 * PVIF ( 9%,20) = $ 208,033 * 0.1784= $ 37,119

4.  1.6%* 15 * 160,000 = $ 38,400

$ 38,400* 8.7556 = $336,215

$ 336,215 * PVIF ( 9%,17)= $ 336,215 * 0.231 = $ 77,690
Answer #2

For number 4. It should be:

 1.6% * 18 * $160,000 = $46,080

$46,080 * 8.7556 = $403,459

$403,459 * PVIF (9%,17) = $403,459 * 0.231 = $93,227